OpenStudy (anonymous):

f(x)= 4x/x^2-25 need f' and f".....I tried and failed,thanks.

OpenStudy (anonymous):

So far I have f'=4x^-2-4/25, f"=-8/x^3

OpenStudy (inkyvoyd):

\(\Huge f(x)= \frac{4x}{x^2}-25\) or \(\Huge f(x)= \frac{4x}{x^2-25}\)

OpenStudy (anonymous):

second one. should have used ( )

OpenStudy (inkyvoyd):

ok. did you just learn the quotient rule?

OpenStudy (anonymous):

yeah. what I got above is wrong

OpenStudy (anonymous):

should I not have simplified?

OpenStudy (inkyvoyd):

so then take u=4x u'=4 v=x^2-25 v'=2x \(\Huge \frac{4(x^2-25)-8x^2}{(x^2-25)^2}=\frac{-12x^2-100}{(x^2-25)^2}\)

OpenStudy (anonymous):

wouldnt it be -4x^2-100?

OpenStudy (inkyvoyd):

yes.

OpenStudy (anonymous):

so no simplifying past this point?

OpenStudy (inkyvoyd):

\(\Huge \frac{4(x^2-25)-8x^2}{(x^2-25)^2}=\frac{-4x^2-100}{(x^2-25)^2}=-4\frac{x^2+25}{(x^2-25)^2}\) that's probably the most you can do.

OpenStudy (inkyvoyd):

I know it's cut off but I just factored out the 4 fo rthe last part. I would suggest you either expand the (x^2-25)^2 or factor it out completely into (x-5)^2*(x+5)^2

OpenStudy (inkyvoyd):

x^2+25 does not factor into the reals.

OpenStudy (anonymous):

so what we are left with for f'= -4 (x^2+25)/(x-5)^2*(x+5)^2

OpenStudy (inkyvoyd):

yeah, make sure you put parenthesis f'= -4 (x^2+25)/((x-5)^2*(x+5)^2)

OpenStudy (inkyvoyd):

for f'' it's just more grinding. Factor out the -4, and then start with u and v u=x^2+25 u'=2x v=(x^2-25)^2 v'=2(x^2-25)(2x) (chain rule) and just a bunch of algebra next